OpenStudy (anonymous):

Find the x-intercept & determine whether each intercept is odd or even multiplicity. f(x)=x^3(x-1)^2(x-2)

OpenStudy (anonymous):

x^3=0 x=0 odd (x-1)^2=0 x-1=0 x=1 even x-2=0 x=2 odd